# How to choose efficiently the size of the Bethe-Salpeter Equation Hamiltonian for accurate exciton calculations on supercells

### v2

@@ Abstract
The Bethe-Salpeter Equation (BSE) is the workhorse method to study excitons in materials. The BSE Hamiltonian size, which depends on how many valence-to-conduction band transitions are considered, needs to be chosen to be sufficiently large to converge excitons’ energies and wavefunctions but should be minimized to make calculations tractable, as BSE calculations are expensive and scale with the number of atoms as $\mathcal{O}(N_{\rm{atoms}}^{6})$. In particular, in the case of supercell (SC) calculations composed of $N_{\rm{rep}}$ replicas of a primitive cell (PC), a natural choice to build this BSE Hamiltonian is to include all transitions derived from PC calculations by zone folding. However, this leads to a very large BSE Hamiltonian, as the number of matrix elements in it is $(N_{k}N_{c}N_{v})^{2}$, where $N_{k}$ is the number of $k$-points and $N_{c(v)}$ is the number of conduction (valence) states used. When creating a SC, the number of $k$-points decreases by a factor $N_{\rm{rep}}$ but both the number of conduction and valence states increase by the same factor, therefore the number of matrix elements in the BSE Hamiltonian increases by a factor $N_{\rm{rep}}^{2}$, making exactly corresponding calculations prohibitive. Here, we provide a workflow to decide how many transitions are necessary to achieve comparable results, based on only PC results. With our method, we show that to converge the first exciton binding energy of a LiF SC composed of 64 PCs, to an energy tolerance of 0.15 eV, we only need 12% of the valence-to-conduction matrix elements that result from zone folding with a minimal set of bands. The procedure in our work helps in evaluating excitonic properties in large SC calculations, such as moiré patterns, charge density waves, defects, self-trapped excitons, polarons, interfaces, and other kinds of disorder.
@@ I Introduction
DFT studies using the SC approximation are already common [17], but GW and BSE calculations are less common due to their high computational cost [4, 20, 21, 22, 23]. The GW method [24] is known to have good performance in reproducing the experimental bandgaps of semiconductors, while BSE [25] is successful in studying excitonic effects. GW/BSE calculations are performed on top of DFT results and are, in general, more computationally demanding than DFT calculations, and their convergence is harder to achieve. Typical plane-wave DFT calculations scale with $\mathcal{O}(N_{\rm{atoms}}^{3})$ [26], while GW/BSE calculations scale up to $\mathcal{O}(N_{\rm{atoms}}^{5})-\mathcal{O}(N_{\rm{atoms}}^{6})$ [27], where $N_{\rm{atoms}}$ is the total number of atoms. There have been advances in parallel efficiency [4] and various approximations and numerical schemes that can be used to handle bigger SCs. To compute the screened Coulomb interaction (the $W$ in the GW approximation), one can make use of analytical models for the dielectric function [28, 29]. The stochastic pseudobands method speeds up GW calculations from $\mathcal{O}(N^{4})$ to $\mathcal{O}(N^{2})$ [30]. Scissor-operators are also a valid option in the case where the GW corrections follow approximately a linear relation with DFT energy levels [31, 27]. Subsampling schemes deal with $\mathbf{G}$-vectors perpendicular to the layers of 2D systems, which greatly reduces the density of $\mathbf{k}$-point sampling the first BZ [32] for both GW and BSE calculations. One can also expand wavefunctions of SC on the basis of wavefunctions of PC, and then calculate SC kernel matrix elements in terms of PC kernel matrix elements [33, 1, 2]. Dai et al. have developed a variational method that combines finite-momentum excitons and phonons to study polaronic excitons without the necessity of SCs [34, 35]. Tight-binding [36, 37, 38, 39] and machine learning [40, 41] approaches to GW/BSE also provide a significant speed-up but need to be parametrized very accurately. Despite this work, new theories and approaches are still necessary to improve the performance of GW/BSE calculations in bigger SCs.
@@
In this paper, we focus on the BSE calculations, and more specifically on the kernel matrix element calculations, which take into account the electron-hole interaction and are typically the most computationally demanding step of our SC GW/BSE calculations [27]. The kernel calculation scales as $\mathcal{O}(N_{\rm{atoms}}^{5})$ and the absorption step, in which the BSE Hamiltonian matrix is diagonalized, scales as $\mathcal{O}(N_{\rm{atoms}}^{6})$, but typically the prefactor makes this step considerably less expensive [27]. The main convergence issue in this kind of calculation is the number of conduction and valence bands to include in the BSE Hamiltonian. The number of bands can be very large, leading to a large number of node-hours and amount of RAM needed to perform the calculations, as well as a large amount of disk space needed to store wavefunctions and GW/BSE results, making BSE calculations on SCs computationally very challenging. We provide an efficient way of choosing the number of valence and conduction for SC based on the results of PC calculations. Within our scheme, we built a BSE Hamiltonian with only 12% of the number of matrix elements that would result from zone folding of a minimal set of PC conduction and valence states. We achieved kernel calculations 8.4 times faster with errors of only 150 meV for the binding energies of the first exciton in SC calculations compared to PC results for rocksalt LiF.
@@
We chose to study LiF as it is an insulator with strong excitonic effects that are well studied theoretically [25]. Due to the strong electron-phonon interaction, polaronic effects are strong as well [42, 43], and it presents self-trapped excitons with polaronic character [34, 35, 10] and color centers [44, 45, 46]. Therefore LiF is an excellent choice for testing new methodologies for BSE supercell calculations. The paper is organized in the following way: in Section II we discuss the zone-folding aspects to be considered in BSE calculations and our approach, in Section III we provide computational details of our calculations, in Section IV we show our approach applied to LiF SC calculations, and in Section V we present our conclusions.
@@ II Theory
Since the concepts of $k$-point sampling and total number of $\bf{G}$-vectors are common for both DFT and GW/BSE calculations, we start discussing them in the case of DFT. Let’s consider a crystal with lattice vectors $\boldsymbol{a_{1}},\boldsymbol{a_{2}}$ and $\boldsymbol{a_{3}}$ and reciprocal lattice vectors $\boldsymbol{b_{1}},\boldsymbol{b_{2}}$ and $\boldsymbol{b_{3}}$ which obey $\boldsymbol{a_{i}}\cdot\boldsymbol{b_{j}}=2\pi\delta_{ij}$ [47]. If a SC has lattice vectors $N_{1}\boldsymbol{a_{1}}$, $N_{2}\boldsymbol{a_{2}}$, and $N_{3}\boldsymbol{a_{3}}$, with $N_{i}\geq 1$ being a natural number, then the primitive reciprocal lattice vectors are given by $\boldsymbol{b_{1}}/N_{1}$, $\boldsymbol{b_{2}}/N_{2}$, $\boldsymbol{b_{3}}/N_{3}$. Therefore, the SC volume increases by a factor $N_{\rm{rep}}=N_{1}N_{2}N_{3}$, while the first Brillouin Zone (BZ) reciprocal volume decreases by this same factor. In the limit of very large SC, just the $\Gamma$ point is sampled [4]. The results of a SC calculation will be identical to results of a PC calculation when the number $N_{k}^{\rm SC}$ of $k$-points in a regular grid for the SC relates to the number $N_{k}^{\rm PC}$ of $k$-points in a regular grid of a PC as $N_{k}^{\rm SC}=N_{k}^{\rm PC}/N_{\rm{rep}}$. For example, a PC calculation of a crystal with $8\times 8\times 8$ $k$-grid corresponds to a $4\times 2\times 1$ $k$-grid in a $2\times 4\times 8$ SC. Given that $N_{k}^{\rm SC}$ and $N_{k}^{\rm PC}$ are integers, this condition cannot always be met, especially if one considers more general SCs in which the lattice vectors are not multiples but linear combinations of the PC ones, e.g. for the conventional and primitive cells of diamond-structure silicon. In such cases, one can make SC results comparable to PC results by making the density of $k$-points in both BZs similar.
@@
Plane-wave based DFT codes expand periodic quantities as $f(r)=\sum_{\bf{G}}f_{\bf{G}}e^{i\bf{G}\cdot r}$ [48]. Those expansions are done for $\bf{G}$-vectors that obey $\hbar^{2}|\mathbf{G}+{k}|^{2}/2m_{e}\leq E_{\rm{cutoff}}$, where $E_{\rm{cutoff}}$ is the cutoff energy. As the $\bf{G}$-vectors are linear combinations of the primitive reciprocal lattice vectors, the density of $\bf{G}$-vectors in reciprocal space increases by a factor $N_{\rm{rep}}$, so SC calculations will have more $\bf{G}$-vectors for a fixed $E_{\rm{cutoff}}$. Plane-wave DFT codes iteratively diagonalize $N_{k}^{\rm PC}$ Kohn-Sham (KS) Hamiltonians, with a time that scales each as $\mathcal{O}(N_{\bf{G}}^{2}\log{N_{\bf{G}}})$. When dealing with a SC, the number of KS Hamiltonians decreases by a factor $N_{\rm{rep}}$ and each diagonalization will be $\sim N^{2}_{\rm{rep}}$ times more expensive, so the whole SC calculation will be $\sim N_{\rm{rep}}^{-1}\times N_{\rm{rep}}^{2}=N_{\rm{rep}}$ times more expensive. However, we will see that the effect is more dramatic in BSE.
@@
DFT calculations are known to underestimate the bandgap of semiconductors and insulators. One solution to this is many-body perturbation theory. The GW approximation has had success in reproducing the electronic bandgap of a wide variety of materials [24, 27, 49]. To take into account excitonic effects, one can solve the Bethe-Salpeter Equation (BSE) [25, 27], given by $H^{\rm{BSE}}|A\rangle=\Omega_{A}|A\rangle$, where $|A\rangle$ is the exciton eigenvector, $\Omega_{A}$ is the exciton energy and $H^{\rm{BSE}}$ is the BSE Hamiltonian. Expressed in the basis of transitions $v\mathbf{k}\to c\mathbf{k}$ (represented by the ket $|cv\mathbf{k}\rangle$), it is given by
@@
Time (s) Nodes Node-hours Notes
@@
Table 1: Timing of LiF GW/BSE calculations on Frontera at the Texas Advanced Computing Center [51] using CPUs. Each CPU node is a Intel 8280 “Cascade Lake” with 56 cores. Each row corresponds to a different step of the GW/BSE workflow in the BerkeleyGW code [27]. The most demanding step for SC is the kernel calculation (electron-hole interactions). GW/BSE calculations were performed with BerkeleyGW version 4.0, compiled with Intel compilers and OpenMPI. The lowest section corresponds to “naive” choices of the number of bands based only on zone folding from $N_{v}=5,N_{c}=10$ and $N_{v}=3,N_{c}=1$ bands of PC calculations. These timings are extrapolated from the SC calculations, proportional to $\left(N_{c}N_{v}N_{k}\right)^{2}$.
One can try to reduce the size of the BSE Hamiltonian by performing a convergence study, although this process may be very expensive, especially for larger SCs. In our method, we estimate the proper number of valence and conduction bands based only on the results of PC calculations. For an exciton $A$, we define the Weighted Average Single-Particle Energies (WASPEs) as
@@
In Fig. 1, we illustrate the basic idea of the method. An exciton is composed of a linear combination of transitions from valence to conduction states and the energy width of the states that compose that exciton is smaller than the whole width of the band [52, 53], when the transitions that compose this exciton are localized in the Brillouin Zone. In this case, when zone folding, this band will be split into several bands, and some of those may not participate in the composition of that exciton. Therefore, one needs to include in the SC BSE Hamiltonian only the relevant zone-folded bands for the excitons of interest. In practice, this means limiting the sums over the exciton coefficients up to the bands that are inside an energy window $[E^{\rm min},E^{\rm max}]$. Therefore, we estimate the exciton energy from PC calculations as

### v3

@@ Abstract
The Bethe-Salpeter Equation (BSE) is the workhorse method to study excitons in materials. The BSE Hamiltonian size, which depends on how many valence-to-conduction band transitions are considered, needs to be chosen to be sufficiently large to converge excitons’ energies and wavefunctions but should be minimized to make calculations tractable, as BSE calculations are expensive and scale with the number of atoms as $\mathcal{O}(N_{\rm{atoms}}^{6})$. In particular, in the case of supercell (SC) calculations composed of $N_{\rm{rep}}$ replicas of a primitive cell (PC), a natural choice to build this BSE Hamiltonian is to include all transitions derived from PC calculations by zone folding. However, this leads to a very large BSE Hamiltonian, as the number of matrix elements in it is $(N_{k}N_{c}N_{v})^{2}$, where $N_{k}$ is the number of $k$-points and $N_{c(v)}$ is the number of conduction (valence) states used. When creating a SC, the number of $k$-points decreases by a factor $N_{\rm{rep}}$ but both the number of conduction and valence states increase by the same factor, therefore the number of matrix elements in the BSE Hamiltonian increases by a factor $N_{\rm{rep}}^{2}$, making exactly corresponding calculations prohibitive. Here, we provide a workflow to decide how many transitions are necessary to achieve comparable results, based on only PC results. With our method, we show that to converge the first exciton binding energy of a LiF SC composed of 64 PCs, to an energy tolerance of 0.15 eV, we only need 12% of the valence-to-conduction matrix elements that result from zone folding with a minimal set of bands. As an example, we use the number of bands from our method to obtain the absorption spectrum of LiF with a Vk-like defect. The procedure in our work helps in evaluating excitonic properties in large SC calculations, such as moiré patterns, charge density waves, defects, self-trapped excitons, polarons, interfaces, and other kinds of disorder and can be used within any code that builds the BSE Hamiltonian.
@@ I Introduction
DFT studies using the SC approximation are already common [17], but GW and BSE calculations are less common due to their high computational cost [4, 20, 21, 22, 23]. The GW method [24] is known to have good performance in reproducing the experimental bandgaps of semiconductors, while BSE [25, 26, 27, 28, 29] is successful in studying excitonic effects. GW/BSE calculations are performed on top of DFT results and are, in general, more computationally demanding than DFT calculations, and their convergence is harder to achieve. Typical plane-wave DFT calculations scale with $\mathcal{O}(N_{\rm{atoms}}^{3})$ [30], while GW/BSE calculations scale up to $\mathcal{O}(N_{\rm{atoms}}^{5})-\mathcal{O}(N_{\rm{atoms}}^{6})$ [31], where $N_{\rm{atoms}}$ is the total number of atoms. There have been advances in parallel efficiency [4] and various approximations and numerical schemes that can be used to handle bigger SCs. To compute the screened Coulomb interaction (the $W$ in the GW approximation), one can make use of analytical models for the dielectric function [32, 33]. The stochastic pseudobands method speeds up GW calculations from $\mathcal{O}(N^{4})$ to $\mathcal{O}(N^{2})$ [34]. Scissor-operators are also a valid option in the case where the GW corrections follow approximately a linear relation with DFT energy levels [35, 31]. Subsampling schemes deal with $\mathbf{G}$-vectors perpendicular to the layers of 2D systems, which greatly reduces the density of $\mathbf{k}$-point sampling the first Brillouin Zone (BZ) [36] for both GW and BSE calculations. One can also expand wavefunctions of SC on the basis of wavefunctions of PC, and then calculate SC kernel matrix elements in terms of PC kernel matrix elements [37, 1, 2]. Dai et al. have developed a variational method that combines finite-momentum excitons and phonons to study polaronic excitons without the necessity of SCs [38, 39]. Tight-binding [40, 41, 42, 43] and machine learning [44, 45] approaches to GW/BSE also provide a significant speed-up but need to be parametrized very accurately. Despite this work, new theories and approaches are still necessary to improve the performance of GW/BSE calculations in bigger SCs.
@@
In this paper, we focus on the BSE calculations, and more specifically on the kernel matrix element calculations, which take into account the electron-hole interaction and are typically the most computationally demanding step of our SC GW/BSE calculations [31]. The kernel calculation scales as $\mathcal{O}(N_{\rm{atoms}}^{5})$ and the absorption step, in which the BSE Hamiltonian matrix is diagonalized, scales as $\mathcal{O}(N_{\rm{atoms}}^{6})$, but typically the prefactor makes this step considerably less expensive [31]. The main convergence issue in this kind of calculation is the number of conduction and valence bands to include in the BSE Hamiltonian. The number of bands can be very large, leading to a large number of node-hours and amount of RAM needed to perform the calculations, as well as a large amount of disk space needed to store wavefunctions and GW/BSE results, making BSE calculations on SCs computationally very challenging. We provide an efficient way of choosing the number of valence and conduction for SC based on the results of PC calculations. Within our scheme, we built a BSE Hamiltonian with only 12% of the number of matrix elements that would result from zone folding of a minimal set of PC conduction and valence states. We achieved kernel calculations 8.4 times faster with errors of only 150 meV for the binding energies of the first exciton in SC calculations compared to PC results for rocksalt LiF. Herein we use the BerkeleyGW code [31], but this method can be easily extended to other codes that build and diagonalize the BSE Hamiltonian [46, 47, 48, 49].
@@
We chose to study LiF as it is an insulator with strong excitonic effects that are well studied theoretically [25]. Due to the strong electron-phonon interaction, polaronic effects are strong as well [50, 51], and it presents self-trapped excitons with polaronic character [38, 39, 10] and color centers [52, 53, 54]. Therefore LiF is an excellent choice for testing new methodologies for BSE supercell calculations, although our method does not make any assumptions about the material, so it can be applied to any material where excitonic effects are present, such as 2D materials [55, 1, 2, 40, 56], carbon nanotubes [57], etc.
In this work we use the term “exciton” for all eigenvectors of the BSE Hamiltonian, regardless if their energies are below or above the bandgap, as often done in literature on BSE [31]. Usually one reserves the term “exciton” for transitions below the bandgap, as experimentally excitons’ signature are sharp peaks below the continuum in optical absorption experiments, which indicates they are bound excitons, and their binding energy is defined as the difference between the excitation energy and the bandgap. From a more analytical point of view, the binding energy may be defined by the action of the kernel operator [25] ($E_{b}=\langle A|K^{\rm{eh}}|A\rangle$), and even transitions with energy above the bandgap can be classified as “bound” excitons if they are composed of higher-energy transitions. For example, recently, an experimental paper studied the dark excitons in CrI3 with energies above the bandgap [58].
The paper is organized in the following way: in Section II we discuss the zone-folding aspects to be considered in BSE calculations and our approach, in Section III we provide computational details of our calculations, in Section IV we show our approach applied to LiF SC calculations, in Section V we calculate the absorption spectra of LiF with a defect similar to a Vk center , and in Section VI we present our conclusions.
@@ II.1 Consideration about $k$ folding
Since the concepts of $k$-point sampling and total number of $\bf{G}$-vectors are common for both DFT and GW/BSE calculations, we start discussing them in the case of DFT. Let’s consider a crystal with lattice vectors $\boldsymbol{a_{1}},\boldsymbol{a_{2}}$ and $\boldsymbol{a_{3}}$ and reciprocal lattice vectors $\boldsymbol{b_{1}},\boldsymbol{b_{2}}$ and $\boldsymbol{b_{3}}$ which obey $\boldsymbol{a_{i}}\cdot\boldsymbol{b_{j}}=2\pi\delta_{ij}$ [59]. If a SC has lattice vectors $N_{1}\boldsymbol{a_{1}}$, $N_{2}\boldsymbol{a_{2}}$, and $N_{3}\boldsymbol{a_{3}}$, with $N_{i}\geq 1$ being a natural number, then the primitive reciprocal lattice vectors are given by $\boldsymbol{b_{1}}/N_{1}$, $\boldsymbol{b_{2}}/N_{2}$, $\boldsymbol{b_{3}}/N_{3}$. Therefore, the SC volume increases by a factor $N_{\rm{rep}}=N_{1}N_{2}N_{3}$, while the first BZ reciprocal volume decreases by this same factor. In the limit of very large SC, just the $\Gamma$ point is sampled [4]. The results of a SC calculation will be identical to results of a PC calculation when the number $N_{k}^{\rm SC}$ of $k$-points in a regular grid for the SC relates to the number $N_{k}^{\rm PC}$ of $k$-points in a regular grid of a PC as $N_{k}^{\rm SC}=N_{k}^{\rm PC}/N_{\rm{rep}}$. For example, a PC calculation of a crystal with $8\times 8\times 8$ $k$-grid corresponds to a $4\times 2\times 1$ $k$-grid in a $2\times 4\times 8$ SC. Given that $N_{k}^{\rm SC}$ and $N_{k}^{\rm PC}$ are integers, this condition cannot always be met, especially if one considers more general SCs in which the lattice vectors are not multiples but linear combinations of the PC ones, e.g. for the conventional and primitive cells of diamond-structure silicon. In such cases, one can make SC results comparable to PC results by making the density of $k$-points in both BZs similar.
@@
Plane-wave based DFT codes expand periodic quantities as $f(r)=\sum_{\bf{G}}f_{\bf{G}}e^{i\bf{G}\cdot r}$ [59]. Those expansions are done for $\bf{G}$-vectors that obey $\hbar^{2}|\mathbf{G}+{k}|^{2}/2m_{e}\leq E_{\rm{cutoff}}$, where $E_{\rm{cutoff}}$ is the cutoff energy. As the $\bf{G}$-vectors are linear combinations of the primitive reciprocal lattice vectors, the density of $\bf{G}$-vectors in reciprocal space increases by a factor $N_{\rm{rep}}$, so SC calculations will have more $\bf{G}$-vectors for a fixed $E_{\rm{cutoff}}$. Plane-wave DFT codes iteratively diagonalize $N_{k}^{\rm PC}$ Kohn-Sham (KS) Hamiltonians, with a time that scales each as $\mathcal{O}(N_{\bf{G}}^{2}\log{N_{\bf{G}}})$. When dealing with a SC, the number of KS Hamiltonians decreases by a factor $N_{\rm{rep}}$ and each diagonalization will be $\sim N^{2}_{\rm{rep}}$ times more expensive, so the whole SC calculation will be $\sim N_{\rm{rep}}^{-1}\times N_{\rm{rep}}^{2}=N_{\rm{rep}}$ times more expensive. This increase can also be thought of as coming from the fact that the PC Hamiltonian is block-diagonal in k-point but the SC Hamiltonian loses this structure when the k-points are folded together. However, we will see that the effect is more dramatic in BSE.
@@
DFT calculations are known to underestimate the bandgap of semiconductors and insulators. One solution to this is many-body perturbation theory. The GW approximation has had success in reproducing the electronic bandgap of a wide variety of materials [24, 31, 60]. To take into account excitonic effects, one can solve the Bethe-Salpeter Equation (BSE) [25, 31, 29], given by $H^{\rm{BSE}}|A\rangle=\Omega_{A}|A\rangle$, where $|A\rangle$ is the exciton eigenvector, $\Omega_{A}$ is the exciton energy and $H^{\rm{BSE}}$ is the BSE Hamiltonian. Expressed in the basis of transitions $v\mathbf{k}\to c\mathbf{k}$ (represented by the ket $|cv\mathbf{k}\rangle$), it is given by
@@
Calculation Step Time (s) Nodes Node-hours Notes
@@
Table 1: Timing of LiF GW/BSE calculations on Frontera at the Texas Advanced Computing Center [62] using CPUs for LiF PC and SC ($\times4$$\times$4). Each CPU node is a Intel 8280 “Cascade Lake” with 56 cores. Each row corresponds to a different step of the GW/BSE workflow in the BerkeleyGW code labeled by the name of the executable [31]. Parabands is the executable used to generate empty states to be used in GW calculations, Epsilon calculates the dielectric matrix, Sigma calculates GW corrections, Kernel calculates electron-hole interactions, and Absorption uses data from previous steps to interpolate, build, and diagonalize the BSE Hamiltonian, obtaining the absorption spectra. More computational details can be found in Section III). The most demanding step for SC is the kernel calculation. GW/BSE calculations were performed with BerkeleyGW version 4.0, compiled with Intel compilers and OpenMPI. The lowest section corresponds to “naive” choices of the number of bands based only on zone folding from $N_{v}=5,N_{c}=10$ and $N_{v}=3,N_{c}=1$ bands of PC calculations. These timings are extrapolated from the SC calculations, proportional to $\left(N_{c}N_{v}N_{k}\right)^{2}$. For GW calculations $n_{\rm{bands}}$ is the number of calculated bands.
One can try to reduce the size of the BSE Hamiltonian by performing a convergence study, although this process may be very expensive, especially for larger SCs. In our method, we estimate the proper number of valence and conduction bands based only on the results of PC calculations.
II.2 Exciton coefficients over the Brillouin Zone
For an exciton $A$, we define the Weighted Average Single-Particle Energies (WASPEs) as
@@ II.2 Exciton coefficients over the Brillouin Zone
In Fig. 1, we illustrate the basic idea of the method. An exciton is composed of a linear combination of transitions from valence to conduction states and the energy width of the states that compose that exciton is smaller than the whole width of the band [63, 56], when the transitions that compose this exciton are localized in the Brillouin Zone. In this case, when zone folding, this band will be split into several bands, and some of those may not participate in the composition of that exciton. Therefore, one needs to include in the SC BSE Hamiltonian only the relevant zone-folded bands for the excitons of interest. In practice, this means limiting the sums over the exciton coefficients up to the bands that are inside an energy window $E^{\rm min},E^{\rm max}]$. We estimate the exciton energy from PC calculations as
@@
The more the coefficients $A_{cv\mathbf{k}}$ are spread in energy, the more bands will be necessary to build the BSE Hamiltonian, and less computational savings for a given accuracy of the exciton energy can be achieved. Therefore our method can be applied for both Wannier excitons (localized in $\mathbf{k}$) and Frenkel excitons (localized in real space), although one can expect to achieve a greater reduction of the BSE Hamiltonian in the Wannier exciton cases, where transitions are localized in the BZ, such as the ones in Ref. [64]. In the specific case of LiF, the lowest exciton transitions are localized at the $\Gamma$, but our approach can be used within other materials like monolayer MoS2, in which the lowest exciton has energy 1.88 eV and is composed by transitions localized at the K point [55]. We highlight that the current method does not make any assumptions about the materials, so it can be applied to any material where the BSE is applicable.